OpenStudy (anonymous):

Megan owes $26.00 for her text messaging in the month of august. if her text messaging plan costs $9 for the first 650 messages and 10cents for each additional text message, how many text messages did she send that month?

OpenStudy (anonymous):

Wait ill start over. You have the function: \[f(x)=\left\{\eqalign{ &9.00\phantom{.};\phantom{.}x\leq650\\ &0.10x-56.00\phantom{.};\phantom{.}x\gt650\\ }\right.\] So we know that \(f(x)=$26.00\) so we are dealing with the second condition. Let us find out this answer: \[\eqalign{ &f(x)=0.10x-56.00 \\ &26.00=0.10x-56.00 \\ &82.00=0.10x \\ &820=x \\ }\] Therefore, she sent 820 texts